A proposed null hypothesis states that there is no difference between the mean weights of the population of two neighboring districts. The sample mean difference is 12 kilograms, and the standard deviation of the distribution of the difference between sample means is 9 kilograms.Which statement is true?

A. The null hypothesis must be accepted if we choose the 68% confidence level.

B. The null hypothesis must be accepted if we choose the 95% confidence level.

C. The null hypothesis must be rejected if we choose the 99.7 % confidence level.

D. The null hypothesis must be rejected if we choose the 100% confidence level.

E. The null hypothesis must be rejected if we choose the 95% confidence level.

(This is all the information that is given)
Answer #1

Using Empirical rule,

68% CI for true mean difference: (mean-SD,mean+SD)=(12-9,12+9)=(3, 21).

95% CI for true mean difference: (mean-2*SD,mean+2*SD)=(12-2*9,12+2*9)=(-6, 30).

Since 95% CI for true mean difference contains zero hence The null hypothesis must be accepted if we choose the 95% confidence level.

Option: B. The null hypothesis must be accepted if we choose the 95% confidence level.